OpenStudy (surana):

Find the solution to x^2 (exponent of 2) = 36.

OpenStudy (alexandervonhumboldt2):

x^2=36 now do the inverse of ^2. the inverse is sqrt. now sqrt36=6 and -6

OpenStudy (alexandervonhumboldt2):

x=sqrt36=6 , -6

OpenStudy (surana):

So, basically, I do the inverse of squaring 36, which was divided into six, but also negative six. I think I see what you mean, so it should be +/- six, or the positive sign with the negative right under it followed by 6, right?

OpenStudy (alexandervonhumboldt2):

well - numebr and + number in ^2 is always positive. no basicily -6^2=36 and 6^2=36 and sqrt36=6 and -6

OpenStudy (alexandervonhumboldt2):

so yeah the anser is +- 6

OpenStudy (surana):

thank you!

OpenStudy (surana):

Thank you for helping me!

OpenStudy (alexandervonhumboldt2):

np